The following information applies to the questions displayed below .]
In 2021, the Westgate Construction Company entered into a contract to construct a road for Santa Clara County for $10,000,000. The road was completed in 2023. Information related to the contract is as follows:     2021 2022 2023 Cost incurred during the year $2,550,000   $4,250,000   $ 1,870,000  Estimated costs to complete as of year-end 5,950,000     1,700,000     0  Billings during the year 2,050,000    4,750,000    3,200,000  Cash collections during the year   1,825,000     4,100,000    4,075,000  

Westgate recognizes revenue over time according to percentage of completion

2-a. In the journal below, complete the necessary journal entries for the year 2021 (credit "Various accounts" for construction costs incurred). 2-b. In the journal below, complete the necessary journal entries for the year 2022 (credit "Various accounts" for construction costs incurred). 2-c. In the journal below, complete the necessary journal entries for the year 2023 (credit "Various accounts" for construction costs incurred).
Complete this question by entering your answers in the tabs below.
In the journal below, complete the necessary journal entries for the year 2021 (credit "Various accounts" for construction costs incurred). (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)
Req 2A Req 2B Req 2C For all three years
Record construction costs.

Record progress billings.

Record cash collections.

Record gross profit (loss).

all of the requirements thank you
Answer #1

Part 2A

Date

general journal

debit

credit

2021

construction in progress

2550000

Various accounts

2550000

2021

accounts receivable

2050000

Bills on construction contract

2050000

2021

cash

1825000

Accounts receivable

1825000

2021

construction in progress

450000

Cost of construction

2550000

Revenue from long-term contracts

3000000

Revenue = (2550000/(2550000+5950000))*10000000 = 3000000

Part 2 B

Date

general journal

debit

credit

2022

construction in progress

4250000

Various accounts

4250000

2022

accounts receivable

4750000

Bills on construction contract

4750000

2022

cash

4100000

Accounts receivable

4100000

2022

construction in progress

750000

Cost of construction

4250000

Revenue from long-term contract

5000000

Revenue = ((2550000+4250000)/( 2550000+4250000+1700000))*10000000 = 8000000-3000000 = 5000000

Part 2 C

Date

general journal

debit

credit

2023

construction in progress

1870000

Various accounts

1870000

2023

Accounts receivable

3200000

Bills on construction contract

3200000

2023

Cash

4075000

Accounts receivable

4075000

2023

Construction in progress

130000

Cost of construction

1870000

Revenue from long-term contract

2000000

Revenue = 10000000-3000000-5000000 = 2000000